The diameter of the base of a cone is 18 feet. The hight of the cone is 12 feet. What is the volume of the cone?
Answer by dabanfield(803)

The formula for the volume of a cone is:
1/3*area of the base*height
The base is a circle with diameter 18 so the area of the base is pi*r^2 = pi*9^2 = 81*pi.
The area volume of the cone then is:
(1/3)*(81*pi)*12 =
324*pi